OpenStudy (anonymous):

You roll 2 dice. What is the probability that the sum of the dice is greater than 6 and 1 die shows a 2? A 6 X 6 table of dice outcomes will help you to answer this question

OpenStudy (anonymous):

\[\left[\begin{matrix} -& 1 & 2 & 3 & 4 & 5 & 6 \\ 1& 2 & 3 & 4 & 5 & 6 & 7 \\ 2& 3 & 4 & 5 & 6 & 7 & 8\\3& 4 & 5 & 6 & 7 & 8 & 9 \\4& 5 & 6 & 7 & 8 & 9 & 10\\5& 6 & 7 & 8 & 9 & 10 & 11\\6& 7 & 8 & 9 & 10 & 11 & 12\end{matrix}\right]\]

OpenStudy (anonymous):

I answered your question...

OpenStudy (anonymous):

the first row is the values for the 1st die, the first column is the values for the 2nd die

OpenStudy (amistre64):

ow my eyes!!

OpenStudy (anonymous):

the inside consists of the sum of dice, so find all values greater than 6, then of those, take the ones with a value of 2 one of the die

OpenStudy (anonymous):

the chart confuses me :(

OpenStudy (curry):

skittles this question has already been answered in my page look aty my page im takeing the same course

OpenStudy (amistre64):

shows a 2 are these; 1 2 3 4 5 6 1 x 2 x x x x x x 3 x 4 x 5 x 6 x now we need to account for those that are greater than 6 1 2 3 4 5 6 1 2 x x 3 4 5 x 6 x if i did it right

OpenStudy (anonymous):

@curry. im doing this for apex to make up for algebra 2 semester 2. @amister. so i just add all them up and i ge the answer?

OpenStudy (anonymous):

The sum of numbers can be either even or odd.... The probability of having a sum as odd is 1/2. Probability of none of the dices showing a four is (5/6)*(5/6) = 25/36 Thus probability of at least one four is 1 - 25/36 = 11/36 Total probability = 1/2 + 11/36 = 29/36

OpenStudy (amistre64):

yep; there are 4 ways to get a sum greater than 6 AND one of them is a 2 4/36 = 1/9 right?

OpenStudy (anonymous):

got it. thx amistre. new questoin coming up

OpenStudy (amistre64):

another way to do this is just create the sample space and weed out the offending options: 11 21 31 41 51 61 12 22 32 42 52 62 13 23 33 43 53 63 14 24 34 44 54 64 15 25 35 45 55 65 16 26 36 46 56 66 delete everything that doesnt have 2 21 12 22 32 42 52 62 23 24 25 26 now delete everything that sums to 6 or less 52 62 25 26 and thats it; 4 left out of 36